# Conditionally valid Probabilistic Conformal Prediction

### v1

@@ Appendix A Additional results and calculations
The prediction set $\bar{\mathcal{C}}_{\alpha}(x)$ relies on the quantile of the distribution $\frac{1}{n}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}$. However, the proof reveals that this prediction set is equivalent to $\mathcal{C}_{\alpha}(x)$.
@@
where the upper bound only holds if $\{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})\}_{k=1}^{n+1}$ are almost surely distinct.
@@
$\mu=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}.$
Since $\alpha\geq(n+1)^{-1}$, the quantile $Q_{1-\alpha}(\mu)$ is the $k_{\alpha}$th order statistic of $V_{1},\ldots,V_{n}$, where
@@ Appendix B Experimental setup and results
B.1 Algorithm structure and features
We detail the construction of the prediction sets given by CP2. The algorithm is divided into two parts: the first part computes the quantile $Q_{1-\alpha}(\mu)$ (Algorithm 1), and the second part constructs the prediction set $\mathcal{C}_{\alpha}(x)$ (Algorithm 2).
Algorithm 1 Quantile Computation
Input: dataset $\{(X_{k},Y_{k})\}_{k\in[n]}$, significance level $\alpha$, confidence set $\mathcal{R}$, conditional distributions $\Pi_{Y\mid X}$ and $\Pi_{Z\mid X}$, function $f_{\tau}$.
for $k=0$ to $n$ do
     Sample $Z_{k}\sim\Pi_{Z\mid X=X_{k}}$
     $(\bar{\tau}_{k},\bar{\lambda}_{k})\leftarrow$ Equations 3 and 2
     Compute $V_{k}=f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})$
Set $k_{\alpha}=\lceil(1-\alpha)(n+1)\rceil$
$V_{(k_{\alpha})}\leftarrow k_{\alpha}$-th smallest value in $\{V_{k}\}_{k\in[n]}\cup\{\infty\}$
Output: $V_{(k_{\alpha})}$.
Algorithm 2 Prediction Set Computation
Input: new data $x$, dataset $\{(X_{k},Y_{k})\}_{k\in[n]}$, significance level $\alpha$, confidence set $\mathcal{R}$, conditional distributions $\Pi_{Y\mid X}$ and $\Pi_{Z\mid X}$, function $f_{\tau}$.
$Q_{1-\alpha}(\mu)\leftarrow$ Algorithm 1
Sample $z\sim\Pi_{Z\mid X=x}$
$\tau_{x,z}\leftarrow$ Equations 3 and 2
Output: $\mathcal{R}_{z}(x;f_{\tau_{x,z}}(Q_{1-\alpha}(\mu)))$.
Choice of $f_{\tau}$.
We present examples of mappings $f_{\tau}$ and their inverses $f_{\tau}^{-1}$ in Table 2. The choice of the mapping $f_{\tau}$ is crucial for the performance of the method, and we investigate their impact in Section 4. For instance, choosing $f_{\tau}(\lambda)=\tau\lambda$ results in conditionally valid prediction sets, as long as $\Pi_{Y\mid X=x}$ accurately estimates the conditional distribution $\textup{P}_{Y\mid X=x}$; see Corollary 3.3.
Adjustment Trivial Linear Exp Tanh Sigmoid
$f_{\tau}(\lambda)$ $\lambda$ $\tau\lambda$ $\exp(\tau\lambda)$ $\tan(\tau\lambda)$ $(1+\exp(-\lambda\tau))^{-1}$
$f_{\tau}^{-1}(\lambda)$ $\lambda$ $\tau^{-1}\lambda$ $\tau^{-1}\log\lambda$ $\tau^{-1}\mathrm{arctan}\,\lambda$ $\tau^{-1}\log((1-\lambda)^{-1}\lambda)$
Table 2: Adjustment Functions $f_{\tau}$ and their inverses $f_{\tau}^{-1}$.
@@ B.2 Details of the experimental setup
We use the Mixture Density Network [4] implementation from CDE [33] Python package  [^1] https://github.com/freelunchtheorem/Conditional_Density_Estimation as a base model for CP, PCP and CP2. The underlying neural network contains two hidden layers of 100 neurons each and was trained for 1000 epochs for each split of the data. Number of components of the Gaussian Mixture was set to 10 for all datasets.
For the CQR [30] and CHR [35] we use the original authors’ implementation  [^2] https://github.com/msesia/chr. The underlying neural network that outputs conditional quantiles consists of two hidden layers with 64 neurons each. Training was performed for 200 epochs for batch size 250.
@@ B.3 Worst-slab coverage
Here we present some additional experiments related to conditional coverage achieved by different methods. We have used Worst Slab Coverage metric, which is sensitive to the set of labs considered during the search. Following [8, 31], recall that a slab is defined as
@@
In Section 4, we presented the results obtained for $(1-\delta)=0.1$. In this case, the considered slabs must contain at least $10\%$ of the data. In Figure 8 we report results obtained for $(1-\delta)=0.4$. We can see that performance improves a lot compared to $\delta=0.1$, and most results become indistinguishable.
Figure 8: Worst-slab coverage on real data, $(1-\delta)=0.4$.
@@ B.4 Extended results of real data experiments
Table 3 we summarize all metrics from our real-world data experiments. For conditional coverage we report worst-slab coverage with $(1-\delta)=0.1$. On six out of nine datasets CP2 method achieves the best result in conditional coverage.
Table 3: Real data experiments: M. Cov. stands for marginal coverage, C. Cov. is worst-slab coverage (here $(1-\delta)=0.4$) and $w_{sd}$ is average total length of the prediction sets, scaled by standard deviation of $Y$. Nominal coverage level is set to $(1-\alpha)=0.9$. For $\Pi_{Y\mid X}$, PCP, CP2-PCP we use the same underlying mixture density network model with $50$ samples. CHR and $\texttt{CQR}(2)$ also share the same base neural network model. We average results of $50$ random data splits. For each dataset, we highlighted the algorithm achieving the closest conditional coverage.
Dataset Metric CP PCP $\Pi_{Y\mid X}$ CP2-PCP-L CP2-PCP-D CHR CQR CQR2
@@
C. Cov. 0.87 0.89 0.81 0.89 0.88 0.87 0.86 0.87
$w_{sd}$ 0.87 0.92 0.78 0.94 0.93 1.32 1.48 1.30
@@ B.5 Other perspective on conditional coverage
Table 4: Summary results of experiments on real data: New version. M. Cov. stands for marginal coverage, C. Cov. is worst-slab coverage (here $(1-\delta)=0.1$) and $w_{sd}$ is average total length of the prediction sets, scaled by standard deviation of $Y$. Nominal coverage level is set to $(1-\alpha)=0.9$. For $\Pi_{Y\mid X}$, PCP, CP2-PCP we use the same underlying mixture density network model with $50$ samples. CHR and $\texttt{CQR}(2)$ also share the same base neural network model. We average results of $50$ random data splits. For each dataset, we highlighted the algorithm achieving the closest conditional coverage.
$\texttt{CP}^{2}$
Dataset Metric $\Pi_{Y\mid X}$ HYB-D HYB-L PCP-D PCP-L CHR CP CQR CQR2 PCP
bike M. Cov. 0.93 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.92 0.90 0.90 0.89 0.89 0.88 0.79 0.90 0.87 0.85
$w_{sd}$ 0.83 0.81 0.77 0.80 0.79 1.94 0.71 2.25 2.31 0.71
bio M. Cov. 0.91 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.91 0.90 0.90 0.90 0.90 0.90 0.88 0.89 0.89 0.89
$w_{sd}$ 1.95 1.91 1.89 1.95 1.97 1.92 2.34 2.13 2.10 1.89
blog M. Cov. 0.91 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.91 0.90 0.89 0.90 0.89 0.87 0.60 0.87 0.86 0.74
$w_{sd}$ 0.72 0.71 0.60 0.71 0.72 0.31 0.60 0.44 0.39 0.30
fb1 M. Cov. 0.93 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.92 0.89 0.88 0.89 0.88 0.87 0.49 0.90 0.87 0.64
$w_{sd}$ 0.58 0.58 0.49 0.59 0.56 0.26 0.47 0.37 0.33 0.28
fb2 M. Cov. 0.93 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.91 0.88 0.87 0.88 0.88 0.88 0.50 0.89 0.89 0.61
$w_{sd}$ 0.65 0.65 0.53 0.65 0.62 0.33 0.53 0.43 0.37 0.32
meps19 M. Cov. 0.89 0.90 0.90 0.90 0.90 0.90 0.90 0.89 0.90 0.90
C. Cov. 0.89 0.90 0.90 0.89 0.90 0.90 0.54 0.88 0.89 0.78
$w_{sd}$ 1.02 1.04 1.06 1.07 1.19 0.76 1.05 1.14 1.19 0.73
meps20 M. Cov. 0.89 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.89 0.89 0.89 0.90 0.90 0.91 0.58 0.88 0.89 0.80
$w_{sd}$ 0.98 1.00 1.02 1.04 1.15 0.77 1.06 1.09 1.17 0.75
meps21 M. Cov. 0.89 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.89 0.90 0.90 0.89 0.89 0.90 0.54 0.89 0.88 0.81
$w_{sd}$ 0.99 1.01 1.04 1.04 1.16 0.79 1.04 1.13 1.21 0.72
temp M. Cov. 0.82 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
C. Cov. 0.81 0.88 0.89 0.89 0.88 0.86 0.87 0.85 0.86 0.89
$w_{sd}$ 0.78 0.93 0.94 0.93 0.96 1.31 0.87 1.48 1.30 0.92
Appendix C Further discussion on the method
In this section, we address the limitations of CP2. This general framework was designed to combine the advantages of both frequentist and Bayesian methods. The CP2 method can create a broad range of methods by combining $f_{\tau}$ and $\Pi_{x}$ alongside existing conformal techniques.
However, it is important to recognize that our method has its own challenges. Specifically, it relies on an estimator for the conditional distribution $\textup{P}_{Y\mid X}$, which may be difficult to obtain in certain real-world situations. Additionally, our approach might not be compatible with all conformal methods, like with binning procedures.
Our theoretical analysis of the conditional validity illustrates the benefit of more adaptive strategies. These findings may prove helpful to improve the empirical performance of conformal prediction in various problems.
We are excited about the potential of our work to contribute to the development of new uncertainty management techniques. Specifically, when it comes to refining confidence regions obtained via Bayesian methods.
In summary, our results showcase the promise of CP2 in improving the reliability of machine learning models. We hope that our research will motivate further progress in this field and inspire researchers to explore the vast possibilities of this approach.

### v2

@@ A.4 Additional results
While the prediction set $\bar{\mathcal{C}}_{\alpha}(x)$ relies on the quantile of the distribution $\frac{1}{n}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}$, its proof reveals that this prediction set is equivalent to $\mathcal{C}_{\alpha}(x)$.
@@
$\mu_{n}=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}.$
Since $\alpha\geq(n+1)^{-1}$, the quantile $Q_{1-\alpha}(\mu_{n})$ is the $k_{\alpha}$th order statistic of $V_{1},\ldots,V_{n}$, where
@@ Appendix B Experimental Setup and Results
Choice of $f_{t}$.
We present examples of mappings $f_{t}$ and their inverses $f_{\tau}^{-1}$ in Table 2. The choice of the mapping $f_{t}$ is crucial for the performance of the method, and we investigate their impact in Section 4. For instance, choosing $f_{\tau}(\lambda)=\tau\lambda$ results in approximately conditionally valid prediction sets, as long as $\Pi_{Y\mid X=x}$ accurately estimates the conditional distribution $\textup{P}_{Y\mid X=x}$; see Theorem 3.2-Theorem 3.3. Initially we also considered other adjustment functions based on exponent, sigmoid and $\tanh$ functions, but they all performed worse than linear and sum. As we show later in 3, these two selected adjustment function perform similarly, showing only marginal differences on some datasets. Designing new adjustment functions is a possible future research direction.
B.1 Highest predictive density (HPD) regions
CP2 with Explicit Conditional Density estimate: CP2-HPD.
Assume that an estimator the conditional density function is known, denoted by $\gamma_{Y\mid X=x}$. The confidence set is defined as $\mathcal{R}(x;t)=\{y\in\mathcal{Y}\colon\gamma_{Y\mid X=x}(y)\geq-t\}$. We omit the variable $z$ from the notation, as we do not consider exogenous randomization in this case. The parameter $\tau_{x}$ is obtained by solving
$\tau_{x}=\operatornamewithlimits{\arg\min}\left\{\tau\in\mathbb{R}\colon\int_{\mathcal{R}(x;\tau)}\gamma_{Y\mid X=x}(y)\,\mathrm{d}y\geq 1-\alpha\right\}.$ (31)
We then compute $\lambda_{x,y}=-\gamma_{Y\mid X=x}(y)$ and derive the prediction set as
$\mathcal{C}_{\alpha}(x)=\left\{{y\in\mathcal{Y}\colon\gamma_{Y\mid X=x}(y)\geq-f_{\tau_{x}}\left({Q_{1-\alpha}\left({\mu_{n}}\right)}\right)}\right\}.$
If we take $f_{\tau}(\lambda)=\lambda$ and $\varphi=1$, the method shares similarity with the CD-split method, proposed in Izbicki et al. (2020). While CD-split uses $\lambda_{x,y}$ as the conformity score, our method uses $f_{\tau_{x}}^{-1}(\lambda_{x,y})$, which incorporates the information from $\tau_{x}$ to modify $\gamma_{Y\mid X=x}(y)$. The CP2-HPD workflow is summarized in Algorithm 2.
Of course, the computation of (31) is in general highly non-trivial. Izbicki et al. (2020) suggested to use binning, therefore approximating the conditional predictive distribution with histograms. The method is restricted to the case where the dimension of $\mathcal{Y}$ the response is small; see Izbicki et al. (2020) for the case of $\mathcal{Y}=\mathbb{R}$. When the dimension becomes larger, then the estimation of HPD is typically based on Monte Carlo methods, thus requiring the introduction of auxiliary variables.
Algorithm 2 CP2-HPD
Input: dataset $\{(X_{k},Y_{k})\}_{k\in[n]}$, significance $\alpha$, conditional density $\gamma_{Y\mid X}$, function $f_{t}$.
// Compute the $(1-\alpha)$-quantile
for $k=1$ to $n$ do
     Set $\bar{\lambda}_{k}=-\gamma_{Y\mid X=X_{k}}(Y_{k})$
     Set $\bar{\tau}_{k}=\tau_{X_{k}}$ as given in (3)
$Q_{1-\alpha}\left({\mu_{n}}\right)\leftarrow\lceil(1-\alpha)(n+1)\rceil$-th smallest value in $\{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})\}_{k\in[n]}\cup\{\infty\}$
// Compute the prediction set for a new point $x\in\mathbb{R}^{d}$
Compute $\tau_{x}$ in (3).
Output: $\mathcal{C}_{\alpha}(x)=\{y\in\mathcal{Y}\colon\gamma_{Y\mid X=x}(y)\geq-f_{\tau_{x}}({\textstyle{Q_{1-\alpha}({\textstyle{\mu_{n}}})}})\}$.
Name $f_{\tau}(\lambda)$ $f_{\tau}^{-1}(\lambda)$ $\varphi$
Linear $\tau\lambda$ $\tau^{-1}\lambda$ 1
Difference $\tau+\lambda$ $\lambda-\tau$ 0
Table 2: Adjustment Functions $f_{t}$, their inverses $f_{\tau}^{-1}$ and $\varphi$ values used in our experiments.
@@ B.2 Details of the experimental setup
We use the Mixture Density Network (Bishop, 1994) implementation from CDE (Rothfuss et al., 2019) Python package [^2] https://github.com/freelunchtheorem/Conditional_Density_Estimation as a base model for CP, PCP and CP2. The underlying neural network contains two hidden layers of 100 neurons each and was trained for 1000 epochs for each split of the data. Number of components of the Gaussian Mixture was set to 10 for all datasets.
For the CQR (Romano et al., 2019) and CHR (Sesia & Romano, 2021) we use the original authors’ implementation [^3] https://github.com/msesia/chr. The underlying neural network that outputs conditional quantiles consists of two hidden layers with 64 neurons each. Training was performed for 200 epochs for batch size 250.
@@ B.3 Worst-slab coverage
Here we present some additional experiments related to conditional coverage achieved by different methods. We have used Worst Slab Coverage metric, which is sensitive to the set of labs considered during the search. Following Cauchois et al. (2020); Romano et al. (2020b), recall that a slab is defined as
@@
In our experiments we follow Romano et al. (2020b) in our implementation of this metric. Namely, we use $25\%$ of the data to find the worst slab and the use the remaining $75\%$ to calculate the final value on this slab. We use $5000$ randomly sampled directions, that are the same for each algorithm and change for each replication.
@@ B.4 Extended results of real data experiments
Table 3 we summarize all metrics from our real-world data experiments. For conditional coverage we report worst-slab coverage with $(1-\delta)=0.1$. On six out of nine datasets CP2 method achieves the best result in conditional coverage. In terms of interval width PCP method produces the narrowest intervals.As we can see, it happens at the expense of conditional coverage: PCP often achieves significantly lower values.
We also present a more detailed view of set size differences between the methods. In the main part we reported average rank of each method in Figure 8. We ranked the algorithms by their projected area at each test point and averaged the ranks. Here we show raw areas of the projections onto each pairs of axes for sgemm_small dataset in table 4. All targets were standardised to zero mean and unit standard deviation so that different projections will be in the same scale. We see that PCP produces smaller set sizes like in one-dimensional case. Quantile-regression based methods have the largest sets, even larger than the fixed-sized sets of CP. Our approach demonstrates only modest increase in prediction set size compared to PCP while achieving sharper conditional coverage.
Table 3: Summary results of experiments on real data. “M. Cov.” stands for marginal coverage, “C. Cov.” is the worst-slab coverage (here $(1-\delta)=0.1$), and $w_{sd}$ is average total length of the prediction sets, scaled by standard deviation of $Y$. Nominal coverage level is set to $(1-\alpha)=0.9$. For $\Pi_{Y\mid X}$, PCP, CP2-PCP we use the same underlying mixture density network model with $50$ samples. CHR and $\texttt{CQR}(2)$ also share the same base neural network model. We average results of $50$ random data splits. For each dataset, we highlighted the algorithm achieving conditional coverage closest to the nominal level.
$\texttt{CP}^{2}$
Dataset Metric CP PCP $\Pi_{Y|X}$ PCP-L PCP-D CHR CQR CQR2
@@
C. Cov. 0.87 0.89 0.81 0.88 0.89 0.86 0.85 0.86
$w_{sd}$ 0.87 0.92 0.78 0.96 0.93 1.31 1.48 1.30
Table 4: Prediction set size comparison for sgemm_small dataset. Rows correspond to different pairs of targets (dataset has 4 targets). For each method the reported value is the mean area of the 2D projection of the prediction set to the corresponding axes pair.
$\texttt{CP}^{2}$
Axes CP PCP $\Pi_{Y|X}$ PCP-L PCP-D CHR CQR CQR2
(0, 1) 2.137 0.435 0.517 0.576 0.560 2.290 2.550 2.436
(0, 2) 2.145 0.435 0.518 0.577 0.561 2.267 2.506 2.358
(0, 3) 2.145 0.436 0.519 0.578 0.561 2.086 2.366 2.172
(1, 2) 2.146 0.435 0.517 0.576 0.560 2.388 2.622 2.546
(1, 3) 2.146 0.435 0.517 0.576 0.560 2.166 2.461 2.314
(2, 3) 2.154 0.436 0.519 0.578 0.562 2.153 2.430 2.255